OpenStudy (el_tucan):

what is the first step in finding the first derivative using logarithmic differentiation for y = (cos x)^(sin x)?

OpenStudy (anonymous):

Isn't you have to move sin(x) like this => y = sin(x)ln(cos(x))?

OpenStudy (amistre64):

almost, the first step is to just log each side

OpenStudy (anonymous):

yes, i know

OpenStudy (anonymous):

\[y=(\cos x)^{\sin x}\] take a logarithm on the both sides

OpenStudy (amistre64):

ln(y) = sin(x) ln(cos(x))

OpenStudy (anonymous):

then take derivative both side sin = cos and cos = -sin

OpenStudy (el_tucan):

right right right! :)

OpenStudy (el_tucan):

just for referance

OpenStudy (amistre64):

i think you dropped a negative on the ln cosx chain

OpenStudy (amistre64):

\[sin(x)~ln(cos(x))\] \[cos(x)~ln(cos(x))+sin(x)~\frac{-sin(x)}{cos(x)}\] \[cos(x)~ln(cos(x))-sin^2(x)~\frac{1}{cos(x)}\]

OpenStudy (amistre64):

or you might have just went another way to get to the same thing :)

OpenStudy (amistre64):

the end result looks fine tho.

OpenStudy (el_tucan):

thanks huh

OpenStudy (el_tucan):

:) i love open study